# DreamControl: Human-Inspired Whole-Body Humanoid Control for Scene Interaction via Guided Diffusion

### v2

@@ III The DreamControl  methodology for constructing humanoid skills
In DreamControl, we take this latter route of first generating motion plans externally through a pre-trained human motion prior. These generated motion plans are then used implicitly during RL training in the reward signal but not explicitly used as goal conditions by the policy, (hence putting the “Dream” in “DreamControl”). In addition to these dense tracking rewards, we also use sparse and verifiable task-specific rewards to explicitly promote task completion. The overall pipeline is summarized in Fig. 2. We now discuss the two stages of (1) trajectory generation from a human motion prior and (2) RL in more detail. More details about the exact parameters used for each task in these 2 stages are in the Appendix. [^1] Appendix link
@@ III-A Stage 1: Generating Reference Trajectory from a Human Motion Prior
This form of spatiotemporal guidance allows us to connect a generated trajectory to its environment (for example, allowing us to specify where a humanoid should sit, how high it should jump, or the location of the object to be manipulated). Being able to control this interaction point is critical, since in the RL simulator we can instantiate an object at some location and then use trajectories that are guaranteed to approach this object, significantly facilitating the RL exploration problem. In DreamControl, we specifically design the form of spatiotemporal guidance for each task. For instance, our pick task involves providing a spatial target for the wrist. See the Appendix1 for more details on guidance control implementations.
@@
Finally, we apply a layer of post-processing to the generated G1 trajectories prior to passing on to RL. Some of the generated trajectories are not dynamically feasible and thus not fit to be used for tracking in Stage 2. We devise task-specific filtering mechanisms based on some heuristics as discussed in detail in the Appendix.1 We also apply task-specific trajectory refinements to avoid unnecessary movements, such as setting all left arm joints to a default value in the Pick task, where only right arm is used. These are also discussed in the Appendix.1
@@
Each reference trajectory is represented as a sequence of target frames, $\alpha_{i}=[\alpha_{i,0},\,\alpha_{i,\Delta t},...,\,\alpha_{i,(L-1)\Delta t}]$, where $\Delta t=0.05s$ is the time step and $L=196$ is trajectory length (hence each trajectory spans $9.8s$). Each frame is represented as $\alpha_{t}=\{p_{t}^{\text{ref,root}},\,\theta_{t}^{\text{ref,root}},\,q_{t}^{\text{ref}},\,s_{t}^{\text{ref,left}},\,s_{t}^{\text{ref,right}}\}$ where $q_{t}^{\text{ref}}\in\mathbb{R}^{27}$ are the reference joint angles, $p_{t}^{\text{ref,root}}\in\mathbb{R}^{3}$ is the position of the root, $\theta_{t}^{\text{ref,root}}\in\mathbb{R}^{4}$ is the orientation of the root in quaternions, and $s_{t}^{\text{ref,left}},s_{t}^{\text{ref,right}}\in\{0,1\}$ are the left and right hand states with 0 denoting open and $1$ denoting closed. These are manually labeled for each task; for example, for the pick-with-right-hand task, we ensure that the right hand closes immediately after time $t_{g}$ and that the left hand stays closed through the duration of the task. Refer to the Appendix1 for more details.
@@
However, we also explore a method handling certain novel tasks that are not well represented by Omnicontrol training distribution (e.g., pulling drawers) by using IK-based optimization on a base trajectory of a person standing idle (or bending down to pull a drawer below waist-level). More details are described in the Appendix.1
@@ III-B Stage 2: RL with Reference Trajectory
where $(\textbf{t}_{t_{g}}^{\text{b,world}},R_{t_{g}}^{\text{b,world}})$ is the pose of the robot body part link in question in world frame (e.g., the right wrist link for the pick-with-right-hand task) and $(\textbf{t}^{\text{o,b}},R^{\text{o,b}})$ is the offset of the object w.r.t the robot body part link where the object should be placed. All specific offsets and body part links used in each task are listed in the Appendix. 1 We randomize the timestamp $t_{g}$, target positions used to generate trajectories, and thus $p_{t_{g}}$, and other characteristics of the object such as mass and friction. We choose randomization hyperparameters to demonstrate a wide range of settings for which we can generate and use reference trajectories to solve tasks that can be scaled with careful engineering. The exact randomization hyperparameters of the environment are reported in the Appendix.1
Action space: Next we define our action and observation spaces. Our simulated robot is a 27-DoF Unitree G1 equipped with two 7-DoF DEX 3-1 hands, one mounted on each wrist (in real world experiments we use Inspire hands). In this work we restrict hand control to discrete open/closed configurations that are fixed per-task (see Appendix1 for task-specific settings, e.g., extending the right index finger for the open configuration during the button-press task). The action space is therefore defined as $a_{t}\in\mathbb{R}^{29}$, where $a_{t}=\{a_{t}^{\text{body}},a_{t}^{\text{left}},a_{t}^{\text{right}}\}$, with $a_{t}^{\text{body}}\in\mathbb{R}^{27}$ denoting the target joint angles for the G1 body, and $a_{t}^{\text{left}},a_{t}^{\text{right}}\in\mathbb{R}$ controlling the left and right hands, respectively. For the hand controls, negative values correspond to an open hand and positive values to a closed hand.
@@
We also add some (3) task-specific rewards to encourage accomplishing the task with high success rates, such as the reward for raising an object above a height for the pick task. These are described for each task in the Appendix.1 The total reward $r_{t}$ is obtained by:
@@
where $w_{r_{i}}\text{ for }i\in\{1,2,...,10\}$ (indexing over our $10$ reward terms) and $w_{r^{\text{task,sparse}}}$ are task-specific weights whose exact values are given in the Appendix.1
Training: We setup our environment and training in IsaacLab [69] that uses IsaacSim simulation. All policies are trained on an NVIDIA RTX A6000 with $48$ GB vRAM using PPO [70]. For each task, we train for $2000$ iterations with $8192$ parallel environments. See Appendix 1 for more details.
@@ IV-B Simulation Results
Table II reports success rates over $1000$ random environments, with each task success criteria defined in the Appendix.1 Our results show that TaskOnly (a) achieves $0\%$ success across all tasks, since relying solely on sparse rewards provides no dense guidance to discover meaningful motions. TaskOnly+ (b) improves performance by adding engineered dense terms, enabling success on simpler tasks like Press Button and Precise Punch, but still fails on tasks requiring coordinated whole-body motion. For example, in Pick from Ground, the robot must crouch in a balanced manner, and in Jump it must first lower its body before springing upward; with only a pelvis-target reward, the policy instead “settles for” merely stretching its knees but is unable to discover how to perform a real jump (see Fig. 3). TrackingOnly (c) performs better overall, but struggles with fine-grained interactive tasks such as Pick. By combining both tracking and task-specific signals, DreamControl achieves robust performance, outperforming all baselines and achieving the best results on $9$ of $11$ tasks.
@@ IV-C Human-ness Comparison
We also evaluate the human-ness (or naturalness) of trajectories generated by our policies compared to TaskOnly+. For tasks where TaskOnly+ achieves non-zero success, we assess how natural the resulting motions appear. First, we report Fréchet inception distance (FID; [71]) scores on the HumanML3D dataset [68], using task-matched ground-truth trajectories obtained via keyword filtering (see Appendix 1). As shown in Table III, DreamControl consistently achieves lower FID values than TaskOnly+, indicating closer alignment with human motions. An exception occurs in the Pick task, which we conjecture is due to a domain gap: human demonstrations typically involve waist-level picking, whereas the shorter G1 robot often performs shoulder-level picks, making its trajectories less comparable to the human dataset. However, other metrics suggest that our method DreamControlproduces more human-like motions consistently in all tasks.
@@ IV-D Sim2Real Deployment
Deployment: To estimate object positions, we leverage an off-the-shelf open-vocabulary object detection model, OWLv2 [72] for 2D localization, then lift to 3D using depth and object-specific offsets (Fig. 2(C)). Due to OWLv2’s inference latency, we detect the object only in the first frame and hold the estimate fixed thereafter. To mitigate errors from this static estimate, we freeze the lower body during interactive tasks (except bimanual pick) and add a penalty on root velocities to ensure the base remains static. Note that this limitation stems from perception bottlenecks rather than our method; in principle, vision-based policies could be trained via student-teacher distillation (e.g., [4]), which we leave for future work. We provide details for sim2real transfer in the Appendix.1
We successfully deployed policies for: Pick (standing), Bimanual Pick (with boxes of varying weights), Press Button (standing), Open Drawer (at different positions), Precise Punch (standing) and Squat (with varying depths). Representative visualizations are shown in Fig. 1 and more videos are available at <project website link> (Currently, they are in supplementary material for review).
@@ IX-A Task-specific prompts
TABLE X: Text prompts for each task deployed on real robot
Task Prompts
Pick “a person stands in place, grabs the cup from side and lifts up”
Press Button “a person stands in place, presses elevator button”
Bimanual Pick “a person raises the toolbox with both hands”
Squat “a person squats in place and stands up”
@@ IX-D Sim2Real
For sim2real trajectory generation and refinement, we follow a similar design to their corresponding sim tasks with some changes in prompt and trajectory refinements that we will discuss in this section. There are $5$ tasks that we put on hardware robot. Specific prompts used for them are listed in Table X. In addition to the standard refinements as made in the sim versions of these tasks (except Squat), the specific spatial control points and refinements for each task are as follows:-

### v3

@@ III The DreamControlmethodology for constructing humanoid skills
In DreamControl, we take this latter route of first generating motion plans externally through a pre-trained human motion prior. These generated motion plans are then used implicitly during RL training in the reward signal but not explicitly used as goal conditions by the policy, (hence putting the “Dream” in “DreamControl”). In addition to these dense tracking rewards, we also use sparse and verifiable task-specific rewards to explicitly promote task completion. The overall pipeline is summarized in Fig. 2. We now discuss the two stages of (1) trajectory generation from a human motion prior and (2) RL in more detail. More details about the exact parameters used for each task in these 2 stages are in the Appendix.
@@ III-A Stage 1: Generating Reference Trajectory from a Human Motion Prior
This form of spatiotemporal guidance allows us to connect a generated trajectory to its environment (for example, allowing us to specify where a humanoid should sit, how high it should jump, or the location of the object to be manipulated). Being able to control this interaction point is critical, since in the RL simulator we can instantiate an object at some location and then use trajectories that are guaranteed to approach this object, significantly facilitating the RL exploration problem. In DreamControl, we specifically design the form of spatiotemporal guidance for each task. For instance, our pick task involves providing a spatial target for the wrist. See the Appendix for more details on guidance control implementations.
@@
Finally, we apply a layer of post-processing to the generated G1 trajectories prior to passing on to RL. Some of the generated trajectories are not dynamically feasible and thus not fit to be used for tracking in Stage 2. We devise task-specific filtering mechanisms based on some heuristics as discussed in detail in the Appendix. We also apply task-specific trajectory refinements to avoid unnecessary movements, such as setting all left arm joints to a default value in the Pick task, where only right arm is used. These are also discussed in the Appendix.
@@
Each reference trajectory is represented as a sequence of target frames, $\alpha_{i}=[\alpha_{i,0},\,\alpha_{i,\Delta t},...,\,\alpha_{i,(L-1)\Delta t}]$, where $\Delta t=0.05s$ is the time step and $L=196$ is trajectory length (hence each trajectory spans $9.8s$). Each frame is represented as $\alpha_{t}=\{p_{t}^{\text{ref,root}},\,\theta_{t}^{\text{ref,root}},\,q_{t}^{\text{ref}},\,s_{t}^{\text{ref,left}},\,s_{t}^{\text{ref,right}}\}$ where $q_{t}^{\text{ref}}\in\mathbb{R}^{27}$ are the reference joint angles, $p_{t}^{\text{ref,root}}\in\mathbb{R}^{3}$ is the position of the root, $\theta_{t}^{\text{ref,root}}\in\mathbb{R}^{4}$ is the orientation of the root in quaternions, and $s_{t}^{\text{ref,left}},s_{t}^{\text{ref,right}}\in\{0,1\}$ are the left and right hand states with 0 denoting open and $1$ denoting closed. These are manually labeled for each task; for example, for the pick-with-right-hand task, we ensure that the right hand closes immediately after time $t_{g}$ and that the left hand stays closed through the duration of the task. Refer to the Appendix for more details.
@@
However, we also explore a method handling certain novel tasks that are not well represented by Omnicontrol training distribution (e.g., pulling drawers) by using IK-based optimization on a base trajectory of a person standing idle (or bending down to pull a drawer below waist-level). More details are described in the Appendix.
@@ III-B Stage 2: RL with Reference Trajectory
where $(\textbf{t}_{t_{g}}^{\text{b,world}},R_{t_{g}}^{\text{b,world}})$ is the pose of the robot body part link in question in world frame (e.g., the right wrist link for the pick-with-right-hand task) and $(\textbf{t}^{\text{o,b}},R^{\text{o,b}})$ is the offset of the object w.r.t the robot body part link where the object should be placed. All specific offsets and body part links used in each task are listed in the Appendix. We randomize the timestamp $t_{g}$, target positions used to generate trajectories, and thus $p_{t_{g}}$, and other characteristics of the object such as mass and friction. We choose randomization hyperparameters to demonstrate a wide range of settings for which we can generate and use reference trajectories to solve tasks that can be scaled with careful engineering. The exact randomization hyperparameters of the environment are reported in the Appendix.
Action space: Next we define our action and observation spaces. Our simulated robot is a 27-DoF Unitree G1 equipped with two 7-DoF DEX 3-1 hands, one mounted on each wrist (in real world experiments we use Inspire hands). In this work we restrict hand control to discrete open/closed configurations that are fixed per-task (see Appendix for task-specific settings, e.g., extending the right index finger for the open configuration during the button-press task). The action space is therefore defined as $a_{t}\in\mathbb{R}^{29}$, where $a_{t}=\{a_{t}^{\text{body}},a_{t}^{\text{left}},a_{t}^{\text{right}}\}$, with $a_{t}^{\text{body}}\in\mathbb{R}^{27}$ denoting the target joint angles for the G1 body, and $a_{t}^{\text{left}},a_{t}^{\text{right}}\in\mathbb{R}$ controlling the left and right hands, respectively. For the hand controls, negative values correspond to an open hand and positive values to a closed hand.
@@
We also add some (3) task-specific rewards to encourage accomplishing the task with high success rates, such as the reward for raising an object above a height for the pick task. These are described for each task in the Appendix. The total reward $r_{t}$ is obtained by:
@@
where $w_{r_{i}}\text{ for }i\in\{1,2,...,10\}$ (indexing over our $10$ reward terms) and $w_{r^{\text{task,sparse}}}$ are task-specific weights whose exact values are given in the Appendix.
Training: We setup our environment and training in IsaacLab [69] that uses IsaacSim simulation. All policies are trained on an NVIDIA RTX A6000 with $48$ GB vRAM using PPO [70]. For each task, we train for $2000$ iterations with $8192$ parallel environments. See Appendix for more details.
@@ IV-B Simulation Results
Table II reports success rates over $1000$ random environments, with each task success criteria defined in the Appendix. Our results show that TaskOnly (a) achieves $0\%$ success across all tasks, since relying solely on sparse rewards provides no dense guidance to discover meaningful motions. TaskOnly+ (b) improves performance by adding engineered dense terms, enabling success on simpler tasks like Press Button and Precise Punch, but still fails on tasks requiring coordinated whole-body motion. For example, in Pick from Ground, the robot must crouch in a balanced manner, and in Jump it must first lower its body before springing upward; with only a pelvis-target reward, the policy instead “settles for” merely stretching its knees but is unable to discover how to perform a real jump (see Fig. 3). TrackingOnly (c) performs better overall, but struggles with fine-grained interactive tasks such as Pick. By combining both tracking and task-specific signals, DreamControl achieves robust performance, outperforming all baselines and achieving the best results on $9$ of $11$ tasks.
@@ IV-C Human-ness Comparison
We also evaluate the human-ness (or naturalness) of trajectories generated by our policies compared to TaskOnly+. For tasks where TaskOnly+ achieves non-zero success, we assess how natural the resulting motions appear. First, we report Fréchet inception distance (FID; [71]) scores on the HumanML3D dataset [68], using task-matched ground-truth trajectories obtained via keyword filtering (see Appendix ). As shown in Table III, DreamControl consistently achieves lower FID values than TaskOnly+, indicating closer alignment with human motions. An exception occurs in the Pick task, which we conjecture is due to a domain gap: human demonstrations typically involve waist-level picking, whereas the shorter G1 robot often performs shoulder-level picks, making its trajectories less comparable to the human dataset. However, other metrics suggest that our method DreamControlproduces more human-like motions consistently in all tasks.
@@ IV-D Sim2Real Deployment
Deployment: To estimate object positions, we leverage an off-the-shelf open-vocabulary object detection model, OWLv2 [72] for 2D localization, then lift to 3D using depth and object-specific offsets (Fig. 2(C)). Due to OWLv2’s inference latency, we detect the object only in the first frame and hold the estimate fixed thereafter. To mitigate errors from this static estimate, we freeze the lower body during interactive tasks (except bimanual pick) and add a penalty on root velocities to ensure the base remains static. Note that this limitation stems from perception bottlenecks rather than our method; in principle, vision-based policies could be trained via student-teacher distillation (e.g., [4]), which we leave for future work. We provide details for sim2real transfer in the Appendix.
We successfully deployed policies for: Pick (standing), Bimanual Pick (with boxes of varying weights), Press Button (standing), Open Drawer (at different positions), Precise Punch (standing) and Squat (with varying depths). Representative visualizations are shown in Fig. 1 and more videos are available at Link.
@@ IX-A Task-specific prompts
TABLE X: Text prompts and slow-down factor for each task deployed on real robot
Task Prompts Slow down factor
Pick “a person stands in place, grabs the cup from side and lifts up” 2.5
Precise Punch “a person performs a single boxing punch with his right hand” 1.5
Press Button “a person stands in place, presses elevator button” 1.5
Bimanual Pick “a person raises the toolbox with both hands” 1
Squat “a person squats in place and stands up” 1
@@ IX-D Sim2Real
For sim2real trajectory generation and refinement, we follow a similar design to their corresponding sim tasks with some changes in prompt and trajectory refinements that we will discuss in this section. There are $5$ tasks that we put on hardware robot. Specific prompts used for them are listed in Table X. The generated trajectories are also slowed down by a certain factor as also reported in Table X. This is to ensure safety of the motion and executed and minimize the sim-real gap by making the motion slower. It was empirically observed that with the same speed, due to some sim-to-real gap, the executed trajectories exhibited less success rates for precise tasks like pick where the object got pushed a little bit and thus failed to be grasped as the position is not updated in the current design. This slowing down was not required for Squat and Bimanual Pick tasks. In addition to the standard refinements as made in the sim versions of these tasks (except Squat), the specific spatial control points and refinements for each task are as follows:-